Keiko pays a semiannual health insurance premium of $1600, and an annual car insurance premium of $650. How much does he need to set aside each month to be able to make these payments?
Answer by Sync(1)

"Keiko pays a semiannual health insurance premium of $1600"

Semiannual means every 6 months (half a year), so we can take that number $1600 and divide it by 6 to get the cost per month. ($1600/6=$266 and 2/3)

"and an annual car insurance premium of $650"

Annual means every 12 months (a year), so we take $650 and divide it by 12 to get the cost per month. ($650/12=$54.1666666 (repeating))

"How much does he need to set aside each month to be able to make these payments? "

Now we have to add each number together. I used the exact values instead of rounding.

$54.16666 (repeating) + $266 and 2/3 = 320.83333 (repeating), or, rounded to the nearest hundredth: $320.83.

Keiko must set aside $320.83 each month in order to make these payments.